wiki:UC20170508

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 5/08/17

Attending: Nancy Chanover, Ben Williams, Mark Klaene, Jon Holtzman, Russet McMillan, Joanne Hughes, John Bally, Remy Indebetouw (for Anne Verbiscer), Scott Anderson, Rachel Kuzio de Naray, Mary Beth Kaiser, Mark Hammergren, Leslie Hebb, Chip Kobulnicky, Bill Ketzeback, Sarah Tuttle, John Tobin (for John Wisniewski), and Bruce Gillespie


User feedback and comments from institutional representatives:

  • Washington - Scott had nothing new to report.
  • Georgia State - Rachel had nothing new to report.
  • Seattle - Joanne had a time-exchange topic to discuss later.
  • NMSU - Jon had nothing new to report.
  • JHU - Mary Beth said that the APO wiki is still problematic - the VPN 8000 port is not working from JHU. Mary Beth and Jon will discuss this offline.
  • Oklahoma - John Tobin had nothing new to report.
  • Colorado - John Bally had nothing new to report.
  • Wyoming - Chip had nothing new to report.
  • Virginia - Remy said that his users were somewhat disappointed with recent lack of instrument availability.
  • NAPG - Leslie also said that her users were somewhat disappointed with recent lack of instrument availability.
  • Adler - Mark mentioned that the Adler lease for telescope time is ending, and he wanted to thank the observatory and its staff for all the excellent support given over the past several years — the project was a great success, with roughly 4000 NEOS observed.

Telescope and Instruments Report:

Mark highlighted the upcoming visit by the UVA instrument specialists next week — TripleSpec will be opened up for repairs. Otherwise, things have been recently pretty routine, with the exception of some minor contamination seen on DIS frames, and the problem with the echelle, discussed separately below. Mark’s more detailed report follows:

3.5-m Telescope and Instruments Highlights, 4/6/17 through 5/4/17

0) Overview

A late winter storm brought some nice wet snow to the mountain but overall fairly productive period with minimal telescope issues. A lot of activity has picked up with Sunspot and NSO’s departure Oct 1. There is also a tremendous amount of management activity taking place with (after SDSS4) AS-4. A hail storm damaged several buildings and vehicles a couple weeks ago and insurance claims are now being processed. Repairs will start soon.

1) Telescope - Telescope was operational with only minimal issues.

2) Instruments

DIS blue and red cameras continue to be operational along with slitviewer. No real measurable change to contamination level.

TripleSpec is warm and ready for a servicing trip by UVA instrument personnel starting May 15th. It will be moved to the lab shortly before this date. Planned servicing is 1 week plus vacuum pumping and cool down, which adds about 2 weeks.

Agile has been operational this period.

Echelle had been operational with no problems. Due to limited vacuum pump availability over the summer we decided to block out time this period to replace the ion pump. This went well but about 30 hrs after cool-down severe contamination appeared that we can only attribute to condensed water vapor. This caused us to warm the camera and try and sublimate the water vapor by using the CCD chip heater. This is now underway.

NIC-FPS is operational.

ARCTIC is operational with no change.

3) ARCSAT - no change and has seen a fair amount of productive use lately.


ARCES status update

According to Bill and Mark, the echelle (aka ARCES) was being cooled after routine servicing when contamination on the CCD was detected. Bill immediately started warming the camera to 0 degrees Centigrade and held it at that temperature for five days in order to evaporate the assumed moisture on the detector. The camera will be re-cooled gradually, and the instrument might be ready for observing tonight. The cause of the contamination is not presently known, and further research is underway. Nancy thanks everyone for their patience and understanding during this unplanned instrument down time.


Q2 3.5-m and ARCSAT scheduling (esp. in light of previous agenda item):

Ben and Russet reported that there is Director’s Discretionary time in late June that may end up being used for payback for the recent lost echelle time. Otherwise, this DD time will be open for requests to use it. Joanne mentioned that she has recently swapped some time for nights on WIYN.


Q3 3.5-m and ARCSAT scheduling:

Ben announced that the Q3 scheduling request has gone out to the institutional schedulers, and we are waiting for the institutional inputs. People need to be mindful of the scheduled shutdown dates, and hopefully everyone will continue to use the new scheduling request forms.


Discussion of process for getting feedback from ARCSAT users:

Nancy said she feels that the ARCSAT feedback channels are largely hit-or-miss — is a new process needed that better yields feedback from the ARCSAT users? There followed an extended discussion, which concluded that it’s not obvious that much more user feedback is needed, but Nancy noted that the Users Committee members can, and should, solicit monthly user feedback from their users for both the 3.5-m and ARCSAT.


Fiber feed project from 3.5-m to APOGEE (and action item from previous meeting):

Nancy noted that the action item from last month was to re-canvas the user community, particularly those new to ARC, on how much interest there is in connecting the 3.5-m telescope to the APOGEE H-band spectrograph at SDSS via fiber optic cables. Leslie said that users need IR spectroscopy (and especially when TripleSpec is down), and that there should be multi-object capabilities; John Tobin said there was interest at Oklahoma in single-target and/or manually plugged multi-fiber-plate science programs; Mary Beth said that JHU astronomers are interested in stellar chemical abundances; John Bally thought that Colorado users would be interested in multi-object IR spectroscopy in young dense clusters; and Leslie voted for studies of eclipsing binaries that exhibit double lines in the IR spectra. All said, Nancy thanked the committee members for this useful feedback, which she will fold into the science cases that were originally written into the proposal white paper some two years ago.


APO presence at upcoming meeting on “Building the Infrastructure for Time-Domain Alert Science in the LSST Era”:

Nancy mentioned that she is planning to attend and present at the subject workshop in Tucson in a couple of weeks — APO was invited to brainstorm with other ground-based facilities on how to best interact with LSST transients and other related time-domain studies. Scott said that Washington would be interested in this workshop and in some kind of transient followup capability at APO. Mary Beth cautioned that transient followup can be very disruptive to regular observing programs if the followups are not limited in scope. Jon said that the question is whether time-domain observations are in ARC’s scientific interests for the telescope, or would such a program be largely for purposes of generating funds. Nancy said that she thought it could be both, but the key is how would it be done most effectively with minimal impact to the regular observation programs on the telescope. Nancy and Ben said that they will continue to explore ideas and scientific motivations — anyone is welcome to share their thoughts, and the discussion on this topic will continue.


ACTION ITEMS:

open from previous meetings: *Users Committee members to reassess level of user interest and scientific potentials for fiber feed system between the 3.5-m telescope and the APOGEE spectrograph. See Instruments for detailed information on this project.

STATUS: Closed, see above.

new from this meeting: None


Next meeting:

The next Users Committee phonecon will be on Monday 12 June at 1030 MDT — the agenda and other materials will be sent to committee members during the preceding week.

Note: See TracWiki for help on using the wiki.